A flywheel of moment inertia 95 kg m2 rotates about its axis with an angular velocity 60 rad/s. Coaxial with the flywheel is a gear-wheel of moment inertia 5.0 kgm2, initially stationary. The two wheels are then joined together.

Calculate the angular velocity of both wheels together after they have been joined.


1
Expert's answer
2021-10-10T18:11:03-0400

"E_1+E_2=E,"

"E_2=0,"

"E_1=\frac{I_1\omega_1^2}2,"

"E=\frac{(I_1+I_2)\omega_2^2}2,\implies"

"\omega_2=\omega_1\sqrt{\frac{I_1}{I_1+I_2}}=58.5~\frac{rad}s."